If an 8kg mass is pulled up an incline at a constant speed ,determine
1)the tension in the string
2)the coefficient of friction between the 8kg mass and the surface (Assume there is no friction between the cord and the pulley )
1
Expert's answer
2021-08-23T12:41:28-0400

Given:

"m=8.0\:\rm kg"

"v=\rm const"

"\theta"


1) the tension in the string

"T=mg\sin\theta-F_{\rm fric}=mg\sin\theta-\mu mg\cos\theta"

2) the coefficient of friction between the mass and the surface

"\mu=\frac{F_{\rm fric}}{N}=\frac{T-mg\sin \theta}{mg\cos\theta}"